According to the elasticity computation of 0.5, how much would popcorn sales fall if the prices increased by:
40%-
30%-
1
Expert's answer
2020-09-29T09:11:15-0400

"E=\frac{percentage.change.in.quantity. demanded}{percentag. change.in.price}"


Case 1

"0.5=\frac{percentage.change.in.quantity. demanded}{40}"


"percentage.change.in.quantity.demanded =\n0.5*40 = 20 percent"


Case 2

"0.5=\frac{percentage.change.in.quantity. demanded}{30}"


"percentage.change.in.quantity.demanded =\n0.5*30 = 15 percent"